OpenStudy (anonymous):

Solve –4(x + 3) = 16

OpenStudy (vishal_kothari):

-7

OpenStudy (anonymous):

expand first: -4x-12=16 isolate x: -4x=16+12 -4x=28 divide by -4: x=-7

OpenStudy (anonymous):

thank you